OpenStudy (anonymous):

I will GIVE MEDAL OR ANYTHING PLEASE HELP ON THIS PROBLEM PLEASEEE!!!! solve the proportion x/9

OpenStudy (idealist10):

x/9=4/x 9*4=x*x 36=x^2 x=6, -6 Answer: 6, -6